OpenStudy (anonymous):

3 + x = 4 is the same as 4 = 3 + x. This is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (isaiah.feynman):

Commutative property.

OpenStudy (anonymous):

why

OpenStudy (isaiah.feynman):

I'm the math God. :D

OpenStudy (anonymous):

tell i need to see why

OpenStudy (isaiah.feynman):

Its fundamental I think...

OpenStudy (isaiah.feynman):

^_^

OpenStudy (anonymous):

ok it's definitely not going to be A and D so it leaves B and C... here's a problem for Associative Property of Addition (6 + 3) + 4.. and here's a problem of Commutative Property of Addition a + b = b + a.. now you tell me which one looks like 3 + x = 4